OpenStudy (anonymous):

Solve the system of equations and choose the correct ordered pair. 5x + 7y = 13 12x - 5y = -56

OpenStudy (anonymous):

\[x=-3,y=4\]

OpenStudy (anonymous):

use elimination, multiply every element in the top equation by 5 and every element in the bottom equation by 7 25x+35y=65 84x-35y=-392 Notice how the coeefficients in front of the y are equal but opposite, add the two equations thus eliminating the y in the process: 109x +0y=-327

OpenStudy (anonymous):

no solve for x, where x=-327/109 = -3

OpenStudy (anonymous):

knowing the value for x you can substitute it into any of the two equations, choose the easiest one (the one with smaller numbers) 5(-3)+7y=13

OpenStudy (anonymous):

7y=13+15=28 y=28/7=4

OpenStudy (anonymous):

I reached the same conclusion using substitution (which maybe takes longer in this particular problem) instead of the ordered pair approach. My work is in the attached file.